Orange Company bought a machine on January 1, 2016. The machine cost $288,000 and had an expected salvage value of $48,000. The life of the machine was estimated to be 5 years. Using straight line depreciation, the book value of the machine at the beginning of the third year would be: A. $192,000 B. $240,000 C. $ 96,000 D. $144,000
Answer #1

Depreciation per year = (book value - salvage value) / life of the assets

= (288000-48000) /5 = $48,000

Book value of the machine at the beginning of the third year = 288000 - 48000*2

= $192,000

Option A. is correct answer.